OpenStudy (anonymous):

Which of the following represent the zeros of f(x) = 6x^3 - 31x^2 + 4x + 5

OpenStudy (anonymous):

Answer choices:

OpenStudy (arabpride):

(x-5)(3x+1)(2x-1)=0 x=5 x=-1/3 x=1/2 5, –one–third, one–half 5 1/3, 1/2

OpenStudy (anonymous):

Thanks so much!

OpenStudy (arabpride):

YOUr welcome!! Glad to help!! *^.^* Did it make sense???

OpenStudy (anonymous):

Most certainly :)

OpenStudy (arabpride):

ALrighty then *^.^* IF u ever need any help, jus link me to the question, nd ill try me best!!